Update General Info Latest Budget Numbers – No Change (FY13 Department Budgets are due July16 – no word on PIP) Labor – Met with Roger – discussed concern over PIP labor (contractors) – Keep an eye on labor changes – impact to schedule! RLS Info – Elmie – where are we? Shutdown – Comments (Mary)? Task Info (highlights) RFQ – Progress being made – CY Tan will talk about this…. Cavity 7 Refurb – Where are we? Absorber work continuing – Most (if not all) parts are in… Linac Modulator Work – High Power Report (J. Steimel – soon) LCW – work is underway and so far looks to be on schedule - next slide Booster Vacuum – PO for blowers has gone out Booster Transformer – We are proceeding to purchase a new transformer ahead of the PIP schedule due to potential failure of East Gallery transformer Linac Transformer – bigger job then expected – FESS/EE continue to investigate
PIP Webpage and Elog http://www-bd.fnal.gov/cgi-mach/machlog.pl?nb=pip&action=view&page=19&scroll=false&load= We should make better use of our elog… – Simple Entries – Pictures – Weekly work summary? – RFQ has it’s own…..keep separate Lots of data!
LCW The LCW is moving along well: 95 LCW system drained during the week of July 2 Booster choke leak repairs and choke tube replacement pretty much finished Replacing BRF LCW valves as they become available during Booster RF Cavity refurb. 55LCW piping has been cut out at the Long 13 absorber site to make room for the new absorber 95 LCW piping work at the Long 13 absorber site is in progress 95 LCW strainers mounted on enclosure walls, need to finish final hook-ups to 95LCW system 400 MeV valves and hoses have been replaced 95LCW return piping to CUB contractor work is ongoing T&M contractor is looking for aluminum welders to do the 95LCW isolation valve and pump install work
High power test of ferrite cores – As of the last update The high power testing fixture was being worked on at low power – Measuring resonant frequency and Q of the structure (loaded and unloaded) The high current PS was in the room, but not yet connected to the buss inside the high power testing fixture The goal was to complete a high power test of an old ferrite core by the end of June John Reid, Kiyomi Seiya, and Dennis Barak
High power test of ferrite cores – New updates The high current PS has been connected to the center buss and has been tested up to 10-kA – This should be enough bias current for our purposes 5-kW RF amp has been installed An old Stackpole core has been tested up to 3.3- kW with no observable high loss effect An old Toshiba core was tested this morning at about 4-kW with no observable high loss effect John Reid, Kiyomi Seiya, and Dennis Barak
High power test of ferrite cores – What’s next? We need more RF power in order to see the high RF loss effect – The current SS amp that is being used can be swapped with one that has twice as many modules Compare the observed high RF loss effect with documentation from the original test that was done decades ago Begin high power testing of the new ferrite cores John Reid, Kiyomi Seiya, and Dennis Barak
B+dB Booster Cogging Fixed with RPOS feedback Changed by dipole corrector Dipole corrector: 0.009[T-m] @ 24.4[A]
Slot 0 /frontend (Alex) VXI interface to the Acnet Loader for DSP and FPGA codes to the board DSP (Craig) Simplify set up parameters for all devices on the board Memory mapping Protocol for reading and writing parameters and curves. FPGA (Kiyomi) Cogging algorithm List up parameters for Acnet and DSP Cable and Setup Cable (Kent) DC offset to the 24 dipoles at long. Setup (Kent, Kiyomi) Signal and scope at LLRF room Plans